OpenStudy (anonymous):

ln(1 − x) + ln(2 − x) = ln(6)

OpenStudy (anonymous):

Write the left side of the equation as a single logarithm, using the law of logs that lo A - log B = log (A/B) So the left side is just log[(1-x)/(2-x)] = ln 6 Now write an exponential function and solve for x.

OpenStudy (anonymous):

@Easyaspi314 This has a plus sign in-between them.

OpenStudy (anonymous):

Sorry...didnt see the plus sign. So we use the Law of Logs, log A + log B = log (AB) so log (1-x) + log (2-x) = log [(1-x)(2-x)].